# Training Neural Networks on Data Sources with Unknown Reliability

### v2

@@ Results
Noise Type Standard IDPA Co-teaching LAP (Ours)
Original Data 82.81 ± 0.79 83.24 ± 0.56 85.12 ± 0.8 83.2 ± 0.83
Random Permute 83.46 ± 0.91 83.28 ± 0.79 85.6 ± 0.15 83.26 ± 1.09
Random Label 65.01 ± 0.65 64.61 ± 0.71 67.18 ± 0.69 71.95 ± 2.94
Time-series data sources with varied and mixed noise. Figure 3(a) compares the results of not using and using LAP for a time-series classification task on normal and abnormal cardiac rhythms with both label and input noise. To evaluate the model on data sources with differing noise levels, for each number of noisy sources we set the noise levels of sources at linear increments between $25\%$ and $100\%$. Additionally, we linearly increase the number of these noisy sources and measure the area under the precision-recall curve (AUC PR) on the test set. These results shows that LAP allows for improved model training on data with sources of varied noise and with multiple noise types. Figure 3(a) also illustrates that in this setting, LAP is robust to increases in the noise within the dataset, as the AUC PR does not significantly degrade, especially when compared to not using LAP. Interestingly, when only two sources contain noise, the total noise rate in the dataset is only $10\%$ (lower than commonly tested in the literature), which has little negative effect on the standard training method, leading to both methods achieving similar results.
LAP on real-world noisy labels. To further demonstrate the effectiveness of LAP in conjunction with other noisy data methods and on real-world noise, we now present the results on a dataset with human labelling of varied noise rates. Figure 3(b) illustrates the results on CIFAR-10N, a noisy human labelled version of CIFAR-10, where we observe a similar pattern to that depicted in Figure 3(a). In both cases, LAP improved performance on noisy data at all noise levels tested, and the standard deviation over the accuracy is reduced. However, we also note that in this case the performance of LAP degraded faster than in Figure 3(a) as the number of noisy sources increased. This could be for various reasons, such as differing noise types, model baselines, and data type. In Appendix A.10 we also show the results of this experiment when using the same model architecture as in Table 1, and with standard training as a baseline, instead of RRL, where we arrive at the same conclusions.
LAP on a noisy natural language task. We also tested LAP on a natural language task, in which the goal is to predict the sentiment from the movie review. The results are presented in Table 3, for which we tested two types of noise. Note the largest difference between the results is seen for “Random Label" noise, in which LAP clearly out-performs the baselines. This is as expected, since for “Random Permute", tokens in a review could still easily be classified with the correct sentiment as the words indicating a positive or negative review (such as “good" or “bad") will still be contained within the text, but in a different location. Additionally, we expect that when the data is not noisy, using LAP should not limit performance, which we see here. It is interesting however, that Co-teaching performs better in the “Original Data" and “Random Permute" settings. We hypothesise that the IMDB dataset contains some noisy labels (through human error), which were split uniformly across data sources; a limitation of our method that we discuss in Section 5.
@@
Additional results. Many further experiments, such as with varied numbers of sources and source sizes (A.8, A.9 and A.11), models sizes (A.10), Imagenet (A.12 and A.13), a regression task (A.14), and the effect on late training performance (A.13.1) can be found in Appendix which provide further intuitions about LAP and strong evidence for its use in a wide variety of settings.
@@ 5 Discussion
This research reveals some interesting future research directions. In our experiments, we studied models of varied capacity (for example: Table 1 and Appendix A.10, and Table 2 and Appendix A.13), however, it is interesting to further study the effects of ill specified models on noisy data techniques, since most methods assume that models attain smaller losses on the non-noisy data points (for example: in our work and Co-teaching [14]), or that logits are reliable (for example: in RRL [16] and IDPA [15]). Additionally, our work focuses on the setting in which knowing the source of a data point provides extra information in learning from noisy data. However, when the noise level is uniform across all sources, knowing the source value does not provide additional information about a data point’s likelihood of being noisy, and so in this particular case LAP would perform as well as standard training. Here, it would be beneficial to apply a noisy data technique in addition to our method, such as with RRL + LAP, studied in the experiments (Table 2, Figure 3(b), and Appendix A.9).
@@
When using new methods, the computational complexity should also be considered. The additional compute required to apply LAP to a single source is $O(S+B)$ and the extra memory cost is $O(S\times H)$ (where $S$ is the number of unique sources in the dataset, $B$ is the batch size, and $H$ is the length of the loss history), as we calculate the source means and standard deviations online. For multiple sources in a batch, the additional compute becomes $O(S\times S_{b}+B)$ (where $S_{b}$ is the number of unique sources in a batch). The extra memory cost in this case is $O(S\times H+S_{b}\times S)$. Although vectorised in the current implementation, $O(S\times S_{b})$ could be performed with $S_{b}$ parallel jobs of $O(S)$, since each source calculation is independent, significantly improving its speed for larger numbers of sources. This is however, already significantly faster than the baselines tested, in which Co-teaching trains two models, IDPA trains a model twice, and RRL applies a $k$-nearest neighbours approach at each epoch.
@@
Importantly, as is the case with all methods designed for training models on noisy data, thought must be given to the underlying cause of noise. For example, in some scenarios noise within a dataset could be attributed to observations on minority groups during data collection, rather than as a result of errors in measurements, labelling, or data transfer. This is why we began with the assumption that all data (if non-noisy) is expected to be generated from the same underlying probability distribution. If data from minority groups are not carefully considered, any technique for learning from noisy data, or standard training could lead to an unexpected predictive bias [40].
@@
Within this work we presented LAP, a method designed for training neural networks on data generated by many data sources with unknown noise levels. In Section 4, we observed that using LAP during training improves model performance when trained on data generated by a mixture of noisy and non-noisy sources, and does not hinder performance when all data sources are free of noise. The results also illustrated that applying the proposed method is beneficial at varied noise levels, when training on multiple sources with differing levels of noise, and with different types of noise. We additionally see through results in Table 1, 2, 3 and Figure 3(a), and 3(b) that our method is applicable across multiple domains with differing tasks. Moreover, within the Appendix we present many further experiments, showing that: (1) The improvements in results translate to large scale datasets (Imagenet), with multiple noise types (Appendix A.13); (2) LAP is robust to overfitting of noisy data, and therefore achieves significantly greater performance during late training (Appendix A.13.1); and (3) LAP continues to outperform the baseline on a regression task (Appendix A.14). Further, our implementation (Appendix A.1) allows this method to be easily applied to any neural network training where data is generated by multiple sources.
In conclusion, we have shown that the proposed method is versatile – providing improved model performance in a variety of noisy data settings and equal performance on non-noisy data, and therefore envision many scenarios in which LAP is relevant and useful.
@@ Acknowledgments
This study is funded by the UK Dementia Research Institute (UKDRI) Care Research and Technology Centre funded by the Medical Research Council (MRC), Alzheimer’s Research UK, Alzheimer’s Society (grant number: UKDRI–7002), and the UKRI Engineering and Physical Sciences Research Council (EPSRC) PROTECT Project (grant number: EP/W031892/1). Infrastructure support for this research was provided by the NIHR Imperial Biomedical Research Centre (BRC) and the UKRI Medical Research Council (MRC). The funders were not involved in the study design, data collection, data analysis or writing the manuscript.
@@ A.1 Code implementation
An implementation of the proposed method, as well as the code to reproduce the main results in this paper is made available  [^2] https://github.com/alexcapstick/unreliable-sources. The experiments in this work were completed in Python 3.11, with all machine learning code written for Pytorch 2.1 [41]. Other requirements to run the experiments are available in the supplementary code. We make our code available under the MIT license.
All datasets tested are publicly available and easily accessible. Additionally, within the supplementary code we use the default Pytorch Dataset objects for CIFAR-10, CIFAR-100, and F-MNIST and provide Pytorch Dataset objects that will automatically download, unzip, and load the data for PTB-XL, CIFAR-10N, IMDB, and California Housing. We additionally provide code to load Tiny-Imagenet and Imagenet from a local directory, since it requires an agreement before accessing. This makes reproduction of the work presented simple to perform.
Baselines were made available by the authors on GitHub:
All experiments were conducted a single A100 (80GB VRAM), and 32GB of RAM. The full research project required more compute than the experiments reported in the paper to design and implement LAP, as well as choose training parameters for the standard training model that were then used for LAP and the baselines. Further information about the compute time required to run each of the main experiments is found in Appendix A.7, and the corresponding section of the Appendix for further experiments. In total, using this hardware, experiments took approximately 15 days to complete.
@@ A.2 Loss assumptions
Within this work, and for many other methods designed for tackling noisy data, we make use of the assumption that during training, neural networks learn non-noisy patterns before fitting to noisy data, and therefore achieve a greater log-likelihood on non-noisy data during early training. This is discussed within Zhang et al. [3], Arpit et al. [4], Han et al. [14], Arazo et al. [27], Yu et al. [28], Shen and Sanghavi [29], Li et al. [16], Wang et al. [15], which present both theoretical and empirical justifications. We also observe these training dynamics in Appendix A.13.1 (specifically, Figure 5), in which the standard training method overfits to noisy data points significantly during the later training stages when compared to our method for learning from noisy data.
A.3 Further details on design decisions
@@
Firstly, we will discuss why the weighted mean and standard deviation is used for comparing sources with each other. This is done to allow for the identification of noisy sources by our method when different data sources might have significantly different noise levels. This is the case in Figure 2, where we have one source producing $100\%$ noise and other sources with $5\%$, and $2.5\%$ noise levels. Here, because the noise in the sources are so varied, the sources with lower noise levels would likely never have an average loss more than the unweighted mean of the loss trajectory plus the threshold created from the leniency multiplied by the unweighted standard deviation (Algorithm 1). By weighting the mean and standard deviation with the calculated temperature, we are able to ensure that all sources with noise level greater than $0\%$ will be discovered in a number of model update steps proportional to their noise level. This is shown in Figure 1, where the weighted mean and standard deviation of the loss moves further to the left, as the weight reduces the influence of the unreliable source (on the right).
@@
When calculating whether a source should be considered more or less unreliable, we need to be able to calculate a reliability compared to the other sources. We therefore use a mean and standard deviation over previous loss values, so that we can calculate a relative reliability. As described in Section 3, it also provides an intuitive idea about the probability that a non-noisy source is incorrectly identified as noisy on a single step (if the means of loss values are assumed to follow a Gaussian distribution, which is not unreasonable according to the central limit theorem).
@@
In Algorithm 1, we clip $C_{s}$ at 00 to ensure that when we have sources with noise levels of different magnitudes, once the sources with larger noise levels have been heavily weighted and so have insignificant effects on the weighted mean and standard deviation, we are able to start reducing the reliability of other noisy sources immediately, rather than waiting for $C_{s}$ to increase from some negative value to 00. Also, since we use the $\tanh^{2}$ function to calculate loss (or gradient) scaling, clipping $C_{s}$ at 00 prevents us from introducing an error in which loss from reliable sources get scaled with the same factor as unreliable sources.
Explanation for using $\boldsymbol{\tanh^{2}}$
Equation 6 describes the scaling of gradients (or equivalently, loss values) to reduce the contributions to model updates from less reliable data sources over time. Although any monotonically increasing function can be used for $f(\cdot):[0,+\infty)\rightarrow[1,+\infty)$, we find that $1/f(C_{s})=1-\tanh^{2}(0.005\cdot\delta\cdot C_{s})$ provides some nice properties: $y=\tanh^{2}x$ has gradient of 00 at $x=0$, and is asymptotic to $y=1$, which are important qualities for calculating the model update scaling. The stationary point at $x=0$ means that small perturbations in the unreliability of each source around $C_{s}=0$ have little effects on the scaling of model updates, which reduces the consequences of randomness in loss values from a given source and ensures that source contributions are only significantly scaled once it is clear that their loss values are consistently larger than those of the other sources. Secondly, since $y=\tanh^{2}x$ is asymptotic to $y=1$, the scale factor, $1-d_{s}$ (Equation 6) for sources that are calculated to be unreliable eventually reduces to 00, allowing the model to “ignore" these data points as if they were masked in late training.
Explanation for scaling with $\boldsymbol{0.005}$
The value of $0.005$ used within the $\tanh^{2}$ function allows for the depression strength parameter $\delta$ to be on the scale of $1$, as shown in Figure 2.
A.4 Alternative interpretation

### v4

@@ 5 Results
Model Types Noise Type Standard IDPA (2021) Co-teaching (2018) CDR (2021) LS (2022a) LAP (Ours) Original Data 80.43 ± 0.05 79.66 ± 0.21 79.7 ± 0.31 80.37 ± 0.1 80.35 ± 0.07 80.44 ± 0.15 Random Label 76.96 ± 0.58 76.05 ± 0.84 77.05 ± 0.7 76.05 ± 0.31 76.02 ± 0.57 78.74 ± 0.41
To supplement the results of PTB-XL, we present LAP on an additional dataset containing real-world data sources, GoEmotions. This dataset allows us to explore the scenario in which source sizes and the classes they contain vary significantly, which could be a common scenario in real-world use cases. As LAP weights sources based on their log-likelihood, it is important to consider the robustness of our method to variations in the distribution of classes in sources, as well as their sizes. In the GoEmotions dataset, our training set contains source sizes in the range of $1$ to $9320$ with a mean size of $1676$ and a standard deviation of $1477$ – providing varied source data distributions (details in Appendix A.15). Table 5 demonstrates the top-5 accuracy of the methods for this task, and suggests that LAP is robust to varied source data distributions in this real-world dataset. Whilst LAP, CDR, Label Smoothing (LS) and standard training perform similar when trained on the original data, using LAP leads to improved top-5 accuracy on the test set when random labels are introduced in training. On GoEmotions, random labelling had small effects on the performance of the standard training method, reducing the top-5 accuracy of the test set from $80.4$ to $77.0$, however, LAP was still able to significantly reduce that performance loss by achieving a top-5 accuracy of the test set of $78.7$. Here, LAP produced a higher top-5 accuracy than the baselines, with less variability in the results. In Appendix A.16 we further test the imbalance in source-class distributions with an extreme example.
Additional results
@@
Many further experiments, such as with varied numbers of sources and source sizes (A.12 and A.14), model sizes (A.13), Imagenet (A.17 and A.18), a regression task (A.19), the effect on late training performance (A.18) and experiments that strain the assumptions of the method (A.16) can be found in the appendix, which provides further intuitions about LAP and strong evidence for its use in a wide variety of settings.
@@ 6 Discussion
This research reveals some interesting future research directions. In our experiments, we study models of varied capacity (for example, Table 1 and Appendix A.13, and Table 2 and Table 4), however, it is interesting to further study the effects of ill-specified models on noisy data techniques, as most methods assume that models attain smaller losses on the non-noisy data points (for example: in our work and Co-teaching [15]), or that logits are reliable (for example: in RRL [17] and IDPA [16]). In addition, we focus on the setting in which knowing the source of a data point provides additional information when learning from noisy data. However, when the noise level is uniform across all sources, the source value does not provide additional information about a data point’s likelihood of being noisy, and so in this case LAP performs as well as standard training. Here, it would be beneficial to apply a noisy data technique in addition to our method, such as with RRL + LAP, studied in the experiments (Table 2, Figure 5(b), and Appendix A.12).
@@
The additional computation required to apply LAP to a single source is $O(S+B)$ and the additional memory cost is $O(S\times H)$ (where $S$ is the number of unique sources in the dataset, $B$ is the batch size and $H$ is the length of the loss history), as we calculate the source means and standard deviations online. For multiple sources in a batch, the additional compute becomes $O(S\times S_{b}+B)$ (where $S_{b}$ is the number of unique sources in a batch). A measured time cost for a simple experiment is given in Appendix A.20. The extra memory cost in this case is $O(S\times H+S_{b}\times S)$. Although vectorised in the current implementation, $O(S\times S_{b})$ could be performed with $S_{b}$ parallel jobs of $O(S)$, since each source calculation is independent, significantly improving its speed for larger numbers of sources. However, this is already significantly faster than some of the baselines tested, in which Co-teaching trains two models, IDPA trains a model for twice as long, and RRL applies $k$-nearest neighbours at each epoch.
@@
Importantly, as is the case with all methods designed for training models on noisy data, consideration must be given to the underlying cause of noise. For example, in some scenarios, noise within a dataset could be attributed to observations on minority groups during data collection, rather than as a result of errors in measurements, labelling, or data transfer. This is why we began with the assumption that all data (if non-noisy) is expected to be generated from the same underlying probability distribution. If data from minority groups are not carefully considered, any technique to learn from noisy data, or standard training could lead to an unexpected predictive bias [44].
7 Conclusion
@@
In this work, we presented LAP, a method designed for training neural networks on data generated by many data sources with unknown noise levels. In Section 5, we observed that using LAP during training improves model performance when trained on data generated by a mixture of noisy and non-noisy sources, and does not hinder performance when all data sources are free of noise. The results also illustrated that applying the proposed method is beneficial at varied noise levels, when training on multiple sources with differing levels of noise, and whilst being robust to different types of noise. We also see through the results in Table 1, 2, 3, 5 and Figure 5(a), and 5(b) that our method is applicable and robust in multiple domains with different tasks. Moreover, within the Appendix we present many further experiments, showing that: (1) Our method is robust to an extreme source class distribution that tests the limits of the assumptions we made when proposing LAP (Appendix A.16); (2) The improvements in results translate to large-scale datasets (Imagenet), with multiple noise types (Appendix 5); (3) LAP is robust to overfitting of noisy data, and therefore achieves significantly greater performance during late training (Appendix A.18); and (4) LAP continues to outperform the baseline on a regression task (Appendix A.19). Furthermore, our implementation (Appendix A.1) allows this method to be easily applied to any neural network training where data are generated by multiple sources, and our analysis in Section 4 and Appendix A.10 provides a detailed description of the parameters introduced and their intuitions.
This work shows that LAP provides improved model performance over the baselines in a variety of noise settings and equal performance on non-noisy data. This is achieved whilst being robust to a multitude of tasks and being cheaper to compute than the baselines designed to tackle noisy data. We therefore imagine many scenarios where LAP is relevant.
Author Contributions
AC: Original idea conception, software, analysis of results, writing, editing, and reviewing; FP: analysis of results, editing, and reviewing; TC: editing, and reviewing; PB: Supervision, analysis of results, reviewing, funding acquisition.
@@ Acknowledgments
This study is funded by the UK Dementia Research Institute (UKDRI) Care Research and Technology Centre funded by the Medical Research Council (MRC), Alzheimer’s Research UK, Alzheimer’s Society (grant number: UKDRI–7002), and the UKRI Engineering and Physical Sciences Research Council (EPSRC) PROTECT Project (grant number: EP/W031892/1). Infrastructure support for this research was provided by the NIHR Imperial Biomedical Research Centre (BRC) and the UKRI Medical Research Council (MRC). The funders were not involved in the study design, data collection, data analysis, or writing the manuscript.
@@ A.1 Code implementation
An implementation of the proposed method as well as the code to reproduce the results in this paper are made available
The experiments in this work were completed in Python 3.11, with all machine learning code written for Pytorch 2.1 [45]. Other requirements for running the experiments are available in the supplementary code. We make our code available under the MIT license.
All tested datasets are publicly available and easily accessible. Additionally, within the supplementary code we use the default Pytorch Dataset objects for CIFAR-10, CIFAR-100, and F-MNIST and provide Pytorch Dataset objects that will automatically download, unzip, and load the data for PTB-XL, CIFAR-10N, IMDB, and California Housing. We additionally provide code to load Tiny-Imagenet and Imagenet from a local directory since it requires an agreement before accessing. This makes reproduction of the work presented simple to perform.
The baselines were made available by the authors on GitHub:
All experiments were carried out on a single A100 (80GB VRAM), and 32GB of RAM. The full research project required more compute than the experiments reported in the paper to design and implement LAP, as well as choose training parameters for the standard training model that were then used for LAP and the baselines. More information on the computation time required to run each of the main experiments is found in Appendix A.9, and the corresponding section of the Appendix for further experiments. In total, using this hardware, experiments took approximately 15 days to complete.
@@ A.2 Loss assumptions
Within this work, and for many other methods designed for tackling noisy data, we make use of the assumption that during training, neural networks learn non-noisy patterns before fitting to noisy data and therefore achieve a greater log-likelihood on non-noisy data during early training. This is discussed in length within the literature, with both theoretical and empirical justifications presented [3, 4, 15, 28, 29, 30, 17, 16, 14]. We also observe these training dynamics in Appendix A.18 (specifically, Figure 9), in which the standard training method overfits to noisy data points significantly during later training stages compared to our method for learning from noisy data.
A.3 Maximum tempered likelihood estimation on clean data
When the dataset contains no noise and batches are large, we can show theoretically that maximising the tempered likelihood provides the same update as the unmodified maximum likelihood training at each parameter step, with a high probability.
We start with the unmodified likelihood, which is iteratively maximised. At every step, because our temperatures are calculated separately from the parameter optimisation, we can assume that they are constant for a single-parameter update. We will denote all data in the data set as $\mathcal{D}$, while $\mathcal{D}_{s}$ refers to all data from the source $s\in S$. Similarly, we will use the shorthand $p_{\theta}(\cdot)=p(\cdot|\theta)$.
$\arg\max_{\theta}\log p_{\theta}(\mathcal{D})$ [Unmodified likelihood]
$=\arg\max_{\theta}\sum_{s\in S}\log p_{\theta}(\mathcal{D}_{s})$ [Observations are independent]
[…]
This provides the parameter $\theta$ that maximises the tempered likelihood. Therefore, if $T_{s}=1$ for all steps, we maximise the same objective as in the unmodified maximum likelihood estimation at each step of the optimisation process.
Since all of the data is non-noisy, the expectation of the log-likelihood of a single observation will remain the same, including across sources. Therefore, given a sample $B$ of observations from $\mathcal{D}$, those observations in $B$ that are contained within each source will have the same mean log-likelihood on average. This is because the expectation of a sample mean is equal to the population mean.
$\mathbb{E}[\frac{1}{|B\cap\mathcal{D}_{i}|}\log p_{\theta}(B\cap\mathcal{D}_{i})]=\mathbb{E}[\frac{1}{|B\cap\mathcal{D}_{j}|}\log p_{\theta}(B\cap\mathcal{D}_{j})]~{}\hskip 10.00002pt\forall i,j\in S$
When the batch size is large, we assume that the observations in $B$ are split by source according to the proportional source sizes. Let $s_{i}=|\mathcal{D}_{s}|/|\mathcal{D}|$ be the proportional size of a source. If the central limit theorem is met, then in a single step, the difference of a given source’s mean log-likelihood ($\mu_{i}=\frac{1}{|B\cap\mathcal{D}_{i}|}\log p_{\theta}(B\cap\mathcal{D}_{i})$) and the mean of all of the other sources’ log-likelihoods ($\mu_{i^{\prime}}=\frac{1}{|B\cap\mathcal{D}_{i}^{\prime}|}\log p_{\theta}(B\cap\mathcal{D}_{i}^{\prime})$) is distributed as:
$\mu_{i^{\prime}}-\mu_{i}\sim N\left(0,\frac{\text{Var}[\log p_{\theta}(\mathcal{D})]}{s_{i}|B|}+\frac{\text{Var}[\log p_{\theta}(\mathcal{D})]}{(1-s_{i})|B|}\right)$
By setting $\text{Var}[\log p_{\theta}(\mathcal{D})]=\sigma_{\text{LL}}^{2}$ and simplifying the fraction, this becomes:
$\mu_{i^{\prime}}-\mu_{i}\sim N\left(0,\frac{\sigma_{\text{LL}}^{2}}{s_{i}(1-s_{i})|B|}\right)$
Therefore, the probability that we increase the perceived unreliability of the source $i$ at a given step with a history length of $1$, using Equation 4 and $\lambda>0$ is:
$p(\text{increase }C_{s})=p\left(\mu_{i^{\prime}}-\mu_{i}>\lambda\sqrt{\frac{\sigma_{\text{LL}}^{2}}{s_{i}(1-s_{i})|B|}}>0\right)>0.5$ (6)
This means that under these assumptions we expect to reduce $C_{s}$ at each step in the temperature calculations, and because we clip $C_{s}$ below at 00, we expect $T_{s}=1$ when all data is clean. When the history length is more than $1$, this can be extended by also including the expected log-likelihood over the different parameters $\theta$ at each model update step.
Equation 6 also theoretically describes how the value of $\lambda$ influences training. This equation can be used to precisely set $\lambda$ to define the probability of incorrectly increasing $C_{s}$ for a given non-noisy source.
Therefore, if the batch size is large enough such that the central limit theorem applies, we expect that the maximum tempered likelihood estimation provides the same value as the maximum likelihood estimation.
A.4 Algorithm details
In the following, we present the implementation of our method as gradient scaling for clarity. Given a dataset $\mathcal{D}$ of features and labels that is generated by sources $S$, $\mathcal{S}=\{s_{1},...,s_{S}\}$. We denote a subset of $\mathcal{D}$ generated by the source $s$ as $\mathcal{D}_{s}\subset\mathcal{D}$, with each data point corresponding to a single source.
In a single update step of a model: we have the perceived unreliability for each source $C=\{C_{s}\}_{1}^{S}$ (initially, all sources are considered reliable and so $C_{s}=0~{}\forall s$); a batch of features, labels, and source values from $\mathcal{D}$; and a history of training losses, $L$ with length $H$ for all sources. Here, $L\in\mathbb{R}^{S\times H}$, and $l_{H}^{s}\in L$ denote the mean loss of data from source $s$ on the most recent batch containing data in $\mathcal{D}_{s}$ (within the subscript $H$). A larger $C_{s}$ denotes a larger estimated noise level for the source $s$. In each step, we update the value $C_{s}$ (and hence the temperature) for a source $s$ using the empirical risk of all other sources $s^{\prime}$ as in Algorithm 1. Further, we define $f(C_{s})$ as:
$f(C_{s})=(1-d_{s})=1-\tanh^{2}(0.005\cdot\delta\cdot C_{s})$
This choice of $f$ has some nice properties discussed in Appendix A.5. We refer to $d_{s}$ as the depression value; and $\delta$ is the depression strength, controlling the rate of depression.
A.5 Further details on design decisions
Intuitive description of Lambda
If no sources are noisy, and the distribution of negative log-likelihoods (NLLs) from the data sources forms a normal distribution, then we incorrectly reduce the temperature of a non-noisy source $s$ with probability $p_{\theta}(\hat{L_{s}}\geq\lambda)$ where $L_{s}=-\log p_{\theta}(\mathcal{D}_{s})$ is the NLL on data from source $s$, and $\hat{L_{s}}$ refers to the standardisation of the NLL using the mean and variance of the NLLs of all other sources $L_{s^{\prime}}$. More details in the influence of a given $\lambda$ value are given in Appendix A.3.
@@ A.5 Further details on design decisions
Firstly, we will discuss why the weighted mean and standard deviation is used for comparing sources with each other. This is done to allow for the identification of noisy sources by our method when different data sources might have significantly different noise levels. This is the case in Figure 4, where we have one source producing $100\%$ noise and other sources with noise levels $5\%$ and $2.5\%$. Here, because the noise in the sources is so varied, sources with lower noise levels would likely never have an average loss more than the unweighted mean of the loss trajectory plus the threshold created from the leniency multiplied by the unweighted standard deviation (Algorithm 1). By weighting the mean and standard deviation with the calculated temperature, we are able to ensure that all sources with a noise level greater than $0\%$ will be discovered in a number of model update steps proportional to their noise level. This is shown in Figure 3, where the weighted mean and standard deviation of the loss moves further to the left, as the weight reduces the influence of the unreliable source (on the right).
@@
When calculating whether a source should be considered more or less unreliable, we need to be able to calculate the reliability compared to the other sources. We therefore use a mean and standard deviation over previous loss values, so that we can calculate a relative reliability. As described in Section 4, it also provides an intuitive idea about the probability that a non-noisy source is incorrectly identified as noisy on a single step (if the means of loss values are assumed to follow a Gaussian distribution, which is not unreasonable according to the central limit theorem).
@@
In Algorithm 1, we clip $C_{s}$ at 00 to ensure that when we have sources with noise levels of different magnitudes, once the sources with larger noise levels have been heavily weighted and therefore have insignificant effects on the weighted mean and standard deviation, we are able to start reducing the reliability of other noisy sources immediately, rather than waiting for $C_{s}$ to increase from some negative value to 00. Also, since we use the $\tanh^{2}$ function to calculate the loss (or gradient) scaling, clipping $C_{s}$ at 00 prevents us from introducing an error in which loss from reliable sources is scaled with the same factor as unreliable sources.
A.6 Alternative interpretation